OpenStudy (anonymous):

Please Please Please help. Solve the system by the substitution method. x2 + y2 = 113 x + y = 15 A. {( -8, 7), ( -7, 8)} B. {( 8, -7), ( 7, -8)} C. {( 8, 7), ( 7, 8)} D. {( -8, -7), ( -7, -8)}

OpenStudy (anonymous):

In the second equation, y = -x - 15 Therefore, x² + (-x-15)² = 113 x² + (-x - 15)(-x - 15) = 113 x² + x² + 30x + 225 = 113 2x² + 30x + 112 = 0 2(x² + 15x + 56) = 0 2(x + 8)(x + 7) = 0 x = -8, -7 If x = -8, then -8 + y = -15 So y = -7 So a solution is (-8, -7) If x = -7, then -7 + y = -15 y = -8 So a solution is (-7, -8) So yeah, the solution is (b).

OpenStudy (anonymous):

i mean d

OpenStudy (anonymous):

Okay thank you so much, you are awesome.